The ratio of men to women at a certain company is 5 to 8 How many employees does the company have if there are 65 male employees?

Found 2 solutions by checkley77, dmf30633:
Answer by checkley77(12569)

65/5=13
13*8=104
13+104=117 TOTAL EMPLOYEES.
5/8=65/104
4*104=8*65
520=520

Answer by dmf30633(3)

The equation you need to formulate would be 5/8=65/x, then solve for x...
5/8=65/x
multiply both sides by 1/65
1/65 * 5/8 = x
5/520 = x
104 = x